A CRV has uniform distribution from -5 to 3. The probability that the CRV is positive is 3/8 A jar contains 3 red marbles, 7 green marbles and 10 white marbles. If a marble is h. drawn from the jar at random, the probability that this marble is white-0.5. Ip 니.lp /- If A and B are two independent events, where PIA 1-0.2 and PIBI-o5 then PIA,B) :0.7 ( F) ip CE) 1p l. The number of different 8-bit words containig three ones and five zeros are 40· 1( lp | The slope of the PDF for CRV couldnt be negative. tǐhiLw k. Two mutually-exclusive events are independent.

True or false with explanation please.
Answer #1

g:

The probability that the CRV is positive is

3 3-(-5) L8J08

Hence, it is true.

h:

Total number of marbles in the jar = 3+7+10 = 20

Out of 20 marbles, 10 are white so

P(white) = 10 /20 = 0.50

Hence, it is true.

i:

False

j:

Since A and B are independent so

P(A and B)=P(A)P(B) = 0.2 *0.5 = 0.10

Hence, it is false.

k:

False

l:

The possible number of permutations of 3 ones and five zeros so

8! 10-56

Hence, it is false.